OpenStudy (anonymous):

The equation h=-16t^2+58t+3 models the height of a baseball t seconds after it has been hit. A. Was the height of the ball ever 40 seconds B. Was the height of the ball ever 60 seconds

OpenStudy (mathstudent55):

Heights are not measured in seconds.

OpenStudy (anonymous):

My bad 40 feet and 60 feet.....

OpenStudy (anonymous):

If I am understanding this correctly, you need to try numbers for the variable "t" and substitute that value in your equation. So, start with 1, which will represent 1 second, do the operation -16(1^2)+.... and so on and see what you get. Is it bigger than 40? if it is then you ball has reached 40 feet at some point . Now, try 2 and 3 and see what happens. At some point you will get smaller heights which means that you should not try more numbers because you have gotten your answer. Physically, your ball in the equation will reach a max height and then after certain time it will come down, hence the smaller numbers that you will obtain after you increase the number of seconds. I hope I was clear.

OpenStudy (bahrom7893):

ok so find out the maximum height of the ball.

OpenStudy (bahrom7893):

h' = -32t + 58 = 0 -32t = -58 t = 58/32 = 1.8125 (units) depends on what the units in the original were. The ball did not rise higher than 1.8125 units

OpenStudy (anonymous):

Thanks much! I gotta go do another page and a half now. Thank you all! Helped me a lot! <3